OpenStudy (confusionist):

A line goes through the following points, and the equation of that line is written in y=mx+b form. What is the equation of the line? xy -10 -43/4 -6 -31/4 -3 -11/2 3 -1 4 -1/4 6 5/4 7 2 How did you find 'b'? It's hardest...

OpenStudy (aripotta):

-10 -43/4 -6 -31/4 -3 -11/2 3 -1 4 -1/4 6 5/4 7 2 these are the points the line goes through?

OpenStudy (confusionist):

Yeah those are them. Its not really the specific problem I need help with, mostly finding b in general.

OpenStudy (aripotta):

well you have to find the slope first. so take two of the points and use the slope formula. i'd go with (3,-1) and (7,2)

OpenStudy (aripotta):

tell me what you get as the slope

OpenStudy (confusionist):

3/4, I believe.

OpenStudy (aripotta):

yup. now we can put the info into point-slope form, then change it into slope-intercept.

OpenStudy (confusionist):

So, ah, we put it into that format how again? I need to be able to do this fast...Right now I can't do it at all lol

OpenStudy (aripotta):

point-slope: y - y1 = m(x - x1) you use one of the points, we'll use (3,-1), and the slope. so y1 will be -1, x1 will be 3, and m will be 3/4

OpenStudy (confusionist):

-1=3/4 x-3 ? I dont see what I would plugin for y1 or x1, so I just plugged in the y term...Or is that confusing?

OpenStudy (aripotta):

y and x stay y and x. y1 is -1, x1 is 3, and m is 3/4 y + 1 = 3/4(x - 3)

OpenStudy (confusionist):

Ah, okay. So my mistake was emitting the variable when I didn't know what went there. What next?

OpenStudy (aripotta):

now we start putting it in slope-intercept form. there's really only 2 steps lol. first distribute 3/4 into the parentheses

OpenStudy (confusionist):

So to do that would we multiply 3 b the three from the parenthesis then divide it by four? Sorry if its way off, my teacher isnt the most intelligent man to walk the earth but he tries.

OpenStudy (aripotta):

multiply 3/4 to x. and yea, divide 3 by 4 and multiply by 3

OpenStudy (aripotta):

so 3 divided by 4 is obviously 3/4. 3/4 times 3 is 2 and 1/4

OpenStudy (aripotta):

does that make sense? now it's y + 1 = 3/4x - 9/4

OpenStudy (confusionist):

Yeah. It's gonna be weird to do this fast on khan academy so I get proficiency. I'd have to do it pretty fast. is b 9/4, then? Oh wait, 1 isnt isolated so we subtract 1 from 4?

OpenStudy (aripotta):

you would subtract 1 from 9/4. or 9/4 - 4/4

OpenStudy (aripotta):

-9/4 - 4/4*

OpenStudy (confusionist):

So when doing 1--9/4 it becomes 1+9/4 because of the double negative? And then you have 12/4? Which simplifies to 1/3? I could be completely wrong, I just went a bit ahead

OpenStudy (aripotta):

nah, just -9 - 4. which is -13. so b is -13/4 or -3 1/4

OpenStudy (aripotta):

...did you understand that? lol

OpenStudy (confusionist):

Yeah took me a minute but Ive got it now lol. Its simpler than I thought lol. Now Im too tired to go on khan and get mastery though lol. Fogettabout homework Im going to bed. So m/slope= 3/4 & b=-13/4?

OpenStudy (aripotta):

yup

OpenStudy (confusionist):

Okay then. Im goin ta sleep. I'll best answer your first reply I suppose. Adios and thanks C:

OpenStudy (aripotta):

have a good sleep :)